Lovie:

A field is to be fertilized at a cost of $0.12 per square yard. The rectangular part of the field is 110 yd long and the diameter of each semicircle is 43 yd. Find the cost of fertilizing the field.

Luigi0210:

Might help if you construct a visual of the problem, do you have one drawn out?

surjithayer:

area of rectangular part=110*43=4730 sq. yards diameter of semicircle=43 yards radius of each semicircle=43 /2 yards area of two semicircles\[=2\left( \frac{ \pi r^2 }{ 2} \right)=\pi r^2=\frac{ 22 }{ 7 }\times \left( \frac{ 43 }{ 2 } \right)^2=1452.79 ~sq.yards\approx ~1453 ~sq.yards\] total area=4730+1453=6183 sq. yards cost of fertilizing the field=6183 *0.12=741.96 $

2O99:

@k1ngofpadlet